OpenStudy (anonymous):

what is the range? y=(1/x+3) +2

OpenStudy (anonymous):

\[y \in \mathbb{R} : y \neq 5\]

OpenStudy (anonymous):

are you sure because the answer choices are (y∈R:y≠2 y∈R:y≠3 y∈R:y≠-2 and y∈R

OpenStudy (anonymous):

I think you typed the problem wrong you said y = (1/x+3) + 2 instead of y = 1/(x+3) + 2 so the range would be: \[y \in \mathbb{R} : y \neq 2\]

OpenStudy (anonymous):

parenthesis just in the wrong place there..

OpenStudy (anonymous):

oh man sorry thanks :)

OpenStudy (anonymous):

np